# Saddle Hierarchy in Dense Associative Memory

### v1

@@ Appendix D Replicated partition function and free entropy
Suppose a student dense associative memory (DAM) model is trained using a dataset $\mathcal{D}=\left\{\mathbf{x}^{c},y^{c}\right\}_{c=1}^{M}$ of $M$ i.i.d. examples $\mathbf{x}^{c}$ with labels $y^{c}$. By Bayes’ theorem, the student weights $\mathbf{w}$ and $\mathbf{p}$ follow the distribution
@@
We will now evaluate the integral $\int_{S^{N-1}}d\mathbf{x}\ I_{y}\left(\mathbf{x}\right)$. Using Eq. (18) of Appendix C with $\rho=\rho_{\gamma_{*}\gamma}:=\frac{1}{N}\sqrt{\sum_{i}\left[\beta^{*}\left[1-\delta_{\gamma_{*}0}\right]w^{*\gamma_{*}}_{i}+\beta\sum_{a}\left[1-\delta_{\gamma_{a}0}\right]w^{a\gamma_{a}}_{i}\right]^{2}}$, we get
@@
Using Eq. (18) of Appendix C, the exponential of $H_{S}$ integrates to
@@
Appendix E Saddle-point equations
@@
In Appendix F, we show that its solution is
@@
for all $0\leq\gamma\leq P$, where $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ is defined using Eqs. (F) of Appendix F and $\sigma_{\gamma}\left(x^{\mu_{*}}\right)=\frac{\exp\left(x^{\mu_{*}\gamma}\right)}{\sum_{\kappa=0}^{P}\exp\left(x^{\mu_{*}\kappa}\right)}$ is the softmax function. We extremize Eq. (22) with respect to the remaining parameters by solving for the gradient equal to zero. $\partial_{m^{\mu_{*}\mu}}f\left(\mathbf{m},\mathbf{\hat{m}},\mathbf{p}\right)=0$ immediately yields
@@
where $\varsigma\left(x\right)=\frac{x}{\sqrt{x^{2}+1}+1}$. $\hat{m}^{\mu_{*}\gamma}$ vanishes (see Eqs. 23) and $m^{\mu_{*}\gamma}$ is arbitrary (see Eqs. 19 and 20) when $\gamma=0$. Therefore, we may update only $\hat{m}^{\mu_{*}\mu}$ and $m^{\mu_{*}\mu}$ with $1\leq\mu\leq P$ when solving for $\nabla f=0$ by fixed-point iteration. Defining $m^{\mu_{*}0}=\frac{1}{\beta_{\text{eff}}}\log\left[\Omega_{N}\left(\beta\right)/\Omega_{N}\left(0\right)\right]$, we then obtain the saddle-point equations
@@
Appendix F Normalization of the weights
$\sum_{\gamma=0}^{P}p^{\gamma}_{y}$ $=p_{\mathbf{q}}\left(y\right)$ (25)
$\sum_{y=0}^{C}p^{\gamma}_{y}$ $=p_{\mathbf{h}}\left(\gamma\right).$
$\frac{\bar{p}^{\gamma}_{y}}{p^{\gamma}_{y}}$ $=\lambda_{y}+\omega^{\gamma},$
$p^{\gamma}_{y}$ $=\frac{1}{\lambda_{y}+\omega^{\gamma}}\bar{p}^{\gamma}_{y}.$
Using the row and column constraints $\sum_{\gamma=0}^{P}p^{\gamma}_{y}=p_{\mathbf{q}}\left(y\right)$ and $\sum_{y=0}^{C}p^{\gamma}_{y}=p_{\mathbf{h}}\left(\gamma\right)$, we find that the Lagrange multipliers $\lambda_{y}$ and $\omega^{\gamma}$ solve the non-linear equations
For conciseness, we define $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)=\lambda_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)+\omega^{\gamma}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$, where $\lambda_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ and $\omega^{\gamma}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ are the $\lambda_{y}$ and $\omega^{\gamma}$ solving Eqs. (F) at given $P$ and $p_{\mathbf{h}}$. Using these definitions, we find the implicit solution $p^{\gamma}_{y}=\bar{p}^{\gamma}_{y}/\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$. As wanted, this equation is useful in analytical calculations involving the saddle-point equations and the stationarity conditions of the loss. However, it is also quite slow to solve numerically for a given $\mathbf{\bar{p}}$, as reported in [109, 110, 111]. Therefore, although we can train $p^{\gamma}_{y}$ by iterating $p^{\gamma}_{y}=\bar{p}^{\gamma}_{y}/\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$, it is not a very efficient method.
In order to efficiently train $p^{\gamma}_{y}$, we devise a faster way to solve Eqs. (F) than through Eq. (F). Exponentiating both sides of the first line of Eqs. (F), we find
where $\eta$ is an arbitrary scalar that will play the role of a learning rate. Fix $k^{\gamma}_{y}=p^{\gamma}_{y}\exp\left(\eta\partial_{p^{\gamma}_{y}}f\left(\mathbf{p}\right)\right)$. By the Sinkhorn-Knopp theorem [112, 113, 114], there is a rescaled matrix of the form $p^{\prime\gamma}_{y}=D^{\gamma}_{L}\left(\mathbf{k}\right)k^{\gamma}_{y}D^{y}_{R}\left(\mathbf{k}\right)$ that satisfies the same constraints $p^{\prime\gamma}_{y}\geq 0$, $\sum_{\gamma=0}^{P}p^{\prime\gamma}_{y}=p_{\mathbf{q}}\left(y\right)$ and $\sum_{y=0}^{C}p^{\prime\gamma}_{y}=p_{\mathbf{h}}\left(\gamma\right)$ as $p^{\gamma}_{y}$ if some technical conditions are satisfied [114]. Moreover, if $p^{\prime\gamma}_{y}$ exists, then it is unique [115, 116, 114], and we can quickly compute suitable scaling factors $D^{\gamma}_{L}\left(\mathbf{k}\right)$ and $D^{\gamma}_{R}\left(\mathbf{k}\right)$ for $\mathbf{k}=\left\{k^{\gamma}_{y}\right\}^{0\leq\gamma\leq P}_{0\leq y\leq C}$ using the Sinkhorn-Knopp algorithm [112, 113, 114, 117]. $p^{\prime\gamma}_{y}$ is generally not equal to $p^{\gamma}_{y}$. However, we observe that the iteration
converges at small $\eta$ (for example $\sim 0.1/P$), which means that it must converge to a solution of Eqs. (F). Since the Sinkhorn-Knopp algorithm is fast, it is significantly more efficient to iterate Eqs. (F) than $p^{\gamma}_{y}=\bar{p}^{\gamma}_{y}/\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ to solve Eqs. (F).
In practice, we train $p^{\gamma}_{y}$ using a stochastic variant of Eqs. (F) where the gradient is estimated over small batches of data and smoothed with a momentum hyperparameter. Furthermore, we multiply the gradient by $\delta_{yy^{\prime}}-p^{\gamma}_{y}/p_{\mathbf{h}}\left(\gamma\right)$ to reduce the size of the components of $\exp\left(\eta\partial_{p^{\gamma}_{y}\left(t\right)}f\left(\mathbf{p}\left(t\right)\right)\right)$ that move $p^{\gamma}_{y}$ away from the constraints $\sum_{\gamma=0}^{P}p^{\gamma}_{y}=p_{\mathbf{q}}\left(y\right)$ and $\sum_{y=0}^{C}p^{\gamma}_{y}=p_{\mathbf{h}}\left(\gamma\right)$. This step is a simple approximation of the gradient projection proposed in [118].
Appendix G Saddle-point hierarchy
@@
where the hidden units $\gamma\in\left\{P+1,...,P+R\right\}$ and their corresponding order parameters are duplicates, or copies, of $\gamma\in\left\{1,...,R\right\}$. $\gamma=0$ can also be duplicated, but the result is less interesting. By definition (see Eqs. F), $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)=\lambda_{y}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)+\omega^{\gamma}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)$, where $\lambda_{y}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)$ and $\omega^{\gamma}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)$ are the $\lambda_{y}$ and $\omega^{\gamma}$ solving
@@
For the rest of this Appendix, all sums are understood as having the same bounds as Eqs. (G) and (3.1). The stability of any fixed point of the form $\mathbf{x}=\mathbf{F}(\mathbf{x})$, such as those of Eqs. (3.1), can be evaluated using the Jacobian matrix $\mathbf{J}$ of $\mathbf{F}$. If all eigenvalues $\lambda$ of the Jacobian satisfy $|\lambda|<1$, then the fixed point is stable. Conversely, if the Jacobian has an eigenvalue $\lambda$ with $|\lambda|>1$, then the fixed point is unstable. In particular, if the quadratic form $\mathbf{v}^{T}\mathbf{J}\mathbf{v}$ is larger than $1$ for some $\mathbf{v}$ with $\|\mathbf{v}\|=1$, then the fixed point is unstable. For the rest of this Appendix, we evaluate the stability of Eqs. (3.1) with duplicated order parameters. Keeping $\mathbf{\bar{p}}$ and $\mathbf{p}$ fixed, the Jacobian of the saddle-point equation for $\mathbf{\bar{x}}$ is
@@
The subcovers $\mathcal{S}\left(0<\mu\leq P\right)$ and $\mathcal{S}\left(R<\mu\leq P+R\right)$ are both partitions of the indices $\mu_{*}$. As such, $\left[\partial_{\bar{x}^{\nu}_{k}}\bar{x}^{\mu}_{j}\right]_{\mu,\nu=1}^{P+R}$ is block-diagonal with $2\times 2$ blocks coupling the indices $0<\mu\leq R$ and $\nu=\mu+P$. Without loss of generality, we investigate the stability of the block
@@
Appendix H Splitting steepest descent
@@
where $\bar{\nabla}_{\bm{\theta}}$ is the gradient constrained to the unit hypersphere $S^{N-1}$ (see Appendix F) and $\beta_{\text{eff}}$ can be written explicitly in terms of $\beta$ as $\beta_{\text{eff}}=\varsigma\beta$ (see Eqs. 12 and 3.1). At step 11 of Alg. (1), we break permutation symmetries in the memories $\mathbf{w}^{\mu}$ by descending along the eigenvectors $\mathbf{u}^{\mu}\in S^{N-1}$ corresponding to the eigenvalues $\lambda_{\text{min}}^{\mu}$. To be more precise, we update the memories $\mathbf{w}^{\mu}$ and their duplicates $\mathbf{w}^{\text{dupli},\mu}$ according to $\mathbf{w}^{\mu}\leftarrow\mathbf{w}^{\mu}+\delta\mathbf{u}^{\mu}$ and $\mathbf{w}^{\text{dupli},\mu}\leftarrow\mathbf{w}^{\text{dupli},\mu}-\delta\mathbf{u}^{\mu}$, respectively, where $\delta$ is a relatively small learning rate [36]. In physics terminology, the eigenvectors $\mathbf{u}^{\mu}$ are excitation modes that break the permutation symmetries of the memories. $N$ and $P$ are generally large, so it is prohibitively expensive to store $\mathcal{S}_{\mu}\left(\mathbf{w},\mathbf{p}\right)$ explicitly for all hidden units $1\leq\mu\leq P$. Therefore, as proposed in [37], we find the eigenvectors $\mathbf{u}^{\mu}$ and their eigenvalues $\lambda_{\text{min}}^{\mu}$ by minimizing the Rayleigh quotients $Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]:S^{N-1}\ni\mathbf{u}^{\mu}\mapsto\left[\mathbf{u}^{\mu}\right]^{T}\mathcal{S}_{\mu}\left(\mathbf{w},\mathbf{p}\right)\mathbf{u}^{\mu}$, which can be evaluated without constructing $\mathcal{S}_{\mu}\left(\mathbf{w},\mathbf{p}\right)$ explicitly. To derive $Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]$, we first compute $\bar{\nabla}_{\mathbf{w}^{\mu}}\bar{\nabla}_{\mathbf{w}^{\mu}}^{T}\exp\left(\beta_{\text{eff}}\sum_{i}w^{\mu}_{i}x_{i}\right)$. Given $f\left(\bm{\theta}\right)=\sum_{i}\theta_{i}x_{i}$, we find
@@
As mentioned at the beginning of Appendix F, $\bar{\nabla}_{\bm{\theta}}f\left(\bm{\theta}\right)$ is the unrestricted gradient of $f\left(\bm{\theta}\right)$ projected onto the tangent space of $\bm{\theta}$. It remains to calculate $\bar{\nabla}_{\bm{\theta}}\bar{\nabla}_{\bm{\theta}}^{T}f\left(\bm{\theta}\right)$. Following [108], we find it to be the iterated projected gradient
@@
This technique is based on the automatic differentiation trick proposed in [37]. We numerically implement the limit in the first equation by setting $\epsilon=0$ during loss evaluation and $\epsilon=1$ during gradient computation. When we optimize $\mathcal{Q}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}\right)$, the minima of the different Rayleigh quotients $Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)$ can span multiple orders of magnitude, so we normalize the gradients $\nabla_{\mathbf{u}^{\mu}}Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)$ by a running average of their magnitudes to facilitate convergence. This is inspired by the use of RMSProp in [37]. Moreover, we constrain the eigenvectors $\mathbf{u}^{\mu}$ and the gradients $\nabla_{\mathbf{u}^{\mu}}Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)$ to the unit hypersphere as we do for $\mathbf{w}^{\mu}$ (see Appendix F). To make Fig. (6), we train the eigenvectors for $1$ epoch, during which we monitor $\min_{\mu}\left\{\frac{1}{2}\left[\mathbf{u}^{\mu}\right]^{T}\nabla_{\mathbf{u}^{\mu}}\mathcal{Q}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)\right\}=\min_{\mu}\left\{\mathcal{Q}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)\right\}\sim\min_{\mu}\left\{\lambda_{\text{min}}^{\mu}\right\}$ as a metric. The Rayleigh quotients converge very quickly, and using a higher patience does not seem to improve the quality of the learned eigenvectors.
Appendix I Initialization and learning rate
To make all our figures, except Fig. (5), we initialize the memories $\mathbf{w}^{\mu}$ uniformly at random on the unit hypersphere $S^{N-1}$ [119]. To make Fig. 5, we instead use the algorithm of [120] to sample the initial memories $\mathbf{w}^{\mu}$ from a vMF distribution (see Appendix C) with mean direction $\tilde{\mathbf{x}}^{*}=\bar{\mathbf{x}}^{*}/\|\bar{\mathbf{x}}^{*}\|$, where $\bar{\mathbf{x}}^{*}=\frac{1}{N}\sum_{\mu_{*}=1}^{P^{*}}\mathbf{x}^{*\mu_{*}}$ is the mean of the patterns $\mathbf{x}^{*\mu_{*}}$ and $\|\bar{\mathbf{x}}^{*}\|=\sqrt{\sum_{i=1}^{N}\left[\bar{x}^{*}_{i}\right]^{2}}$. By construction, the mean direction $\tilde{\mathbf{x}}^{*}$ is the ordered solution of Eq. (3.1) with the most permutation symmetries, or in other words the root of the learning dynamics tree shown in Fig. 5, so initializing the memories around it helps reveal the tree structure.
We use a learning rate of $\eta\sim 0.1$ to train the memories $\mathbf{w}^{\mu}$. Based on our experience, $\mathbf{p}$ trains well when its own learning rate is approximately $\left(1+\frac{p_{\mathbf{h}}\left(0\right)}{1-p_{\mathbf{h}}\left(0\right)}\right)P$ times smaller. Without this rescaling, the multiplicative update factor $\exp\left(\eta\partial_{p^{\gamma}_{y}}L\left(\mathbf{w},\mathbf{p}\right)\right)$ described in Appendix F can be relatively large even when $\eta$ is relatively small, making it ill-behaved. Equivalently, we can also train $g^{\gamma}_{y}=\left(1+\frac{p_{\mathbf{h}}\left(0\right)}{1-p_{\mathbf{h}}\left(0\right)}\right)P\mathbf{p}$ with the same learning rate $\eta$ as $\mathbf{w}$. We adopt this approach in our code available at [38].
[…]
Algorithm 2 Rescaled splitting steepest descent [36, 37]
3:min $L\left(\mathbf{w},\mathbf{g}\right)$ with SGD
4:while $P_{\text{cur}}<P_{\text{max}}$ do
11:  Update $P_{\text{cur}}\leftarrow P_{\text{cur}}+R$
13:  min $L\left(\mathbf{w},\mathbf{g}\right)$ with SGD
14:end while
15:return $\triangleright$ See Appendix H and [36, 37] for details about steps 5 and 12
When we train our DAM without splitting steepest descent, we set $P_{0}=1$. The resulting weights are, in some sense, the most “typical” [121, 122] for the constraints $g^{\gamma}_{y}\geq 0$, $\sum_{\gamma=0}^{P}g^{\gamma}_{y}=\left(P+1\right)p_{\mathbf{q}}\left(y\right)$ and $\sum_{y=0}^{C}g^{\gamma}_{y}=\left(P+1\right)p_{\mathbf{h}}\left(\gamma\right)=1$ (see Section 2). With splitting steepest descent, we set $P_{0}=P_{\text{cur}}/P_{\text{final}}$ to ensure that $g_{\mathbf{h}}\left(0\right)=\sum_{y=0}^{C}g^{0}_{y}$ is approximately the same size as the other entries of $g_{\mathbf{h}}\left(\gamma\right)$ after training. For simplicity, we set $p_{\mathbf{q}}\left(y\right)$ to the proportions of classes in the data during supervised training. For unsupervised training (see Section 4.2), we break the permutation symmetry between the different classes by using different values for all the entries of $p_{\mathbf{q}}\left(y\right)$. Otherwise, class weights with unbroken permutation symmetries stay stuck at their initial conditions.
Appendix J Weights learned with unsupervised training

### v2

@@ Appendix G Replicated partition function and free entropy
Suppose a student DAM model is trained using a dataset $\mathcal{D}=\left\{\mathbf{x}^{c},y^{c}\right\}_{c=1}^{M}$ of $M$ i.i.d. examples $\mathbf{x}^{c}$ with labels $y^{c}$. By Bayes’ theorem, the student weights $\mathbf{w}$ and $\mathbf{p}$ follow the distribution
@@
We will now evaluate the integral $\int_{S^{N-1}}d\mathbf{x}\ I_{y}\left(\mathbf{x}\right)$. Using Eq. (18) of Appendix B with $\rho=\rho_{\gamma_{*}\gamma}:=\frac{1}{N}\sqrt{\sum_{i}\left[\beta^{*}\left[1-\delta_{\gamma_{*}0}\right]w^{*\gamma_{*}}_{i}+\beta\sum_{a}\left[1-\delta_{\gamma_{a}0}\right]w^{a\gamma_{a}}_{i}\right]^{2}}$, we get
@@
Using Eq. (18) of Appendix B, the exponential of $H_{S}$ integrates to
@@
Appendix H Saddle-point equations
@@
In Appendix C, we show that its solution is
@@
for all $0\leq\gamma\leq P$, where $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ is defined using Eqs. (C) of Appendix C and $\sigma_{\gamma}\left(x^{\mu_{*}}\right)=\frac{\exp\left(x^{\mu_{*}\gamma}\right)}{\sum_{\kappa=0}^{P}\exp\left(x^{\mu_{*}\kappa}\right)}$ is the softmax function. We extremize Eq. (26) with respect to the remaining parameters by solving for the gradient equal to zero. $\partial_{m^{\mu_{*}\mu}}f\left(\mathbf{m},\mathbf{\hat{m}},\mathbf{p}\right)=0$ immediately yields
@@
where $\varsigma\left(x\right)=\frac{x}{\sqrt{x^{2}+1}+1}$. $\hat{m}^{\mu_{*}\gamma}$ vanishes (see Eqs. 27) and $m^{\mu_{*}\gamma}$ is arbitrary (see Eqs. 23 and 24) when $\gamma=0$. Therefore, we may update only $\hat{m}^{\mu_{*}\mu}$ and $m^{\mu_{*}\mu}$ with $1\leq\mu\leq P$ when solving for $\nabla f=0$ by fixed-point iteration. Defining $m^{\mu_{*}0}=\frac{1}{\beta_{\text{eff}}}\log\left[\Omega_{N}\left(\beta\right)/\Omega_{N}\left(0\right)\right]$, we then obtain the saddle-point equations
@@
Appendix I Saddle-point hierarchy
@@
where the hidden units $\gamma\in\left\{P+1,...,P+R\right\}$ and their corresponding order parameters are duplicates, or copies, of $\gamma\in\left\{1,...,R\right\}$. $\gamma=0$ can also be duplicated, but the result is less interesting. By definition (see Eqs. C), $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)=\lambda_{y}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)+\omega^{\gamma}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)$, where $\lambda_{y}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)$ and $\omega^{\gamma}\left(\mathbf{\bar{p}}^{\text{dupli}};p_{\mathbf{h}}\right)$ are the $\lambda_{y}$ and $\omega^{\gamma}$ solving
@@
For the rest of this Appendix, all sums are understood as having the same bounds as Eqs. (I) and (3.1). The stability of any fixed point of the form $\mathbf{x}=\mathbf{F}(\mathbf{x})$, such as those of Eqs. (3.1), can be evaluated using the Jacobian matrix $\mathbf{J}$ of $\mathbf{F}$. If all eigenvalues $\lambda$ of the Jacobian satisfy $|\lambda|<1$, then the fixed point is stable. Conversely, if the Jacobian has an eigenvalue $\lambda$ with $|\lambda|>1$, then the fixed point is unstable. In particular, if the quadratic form $\mathbf{v}^{T}\mathbf{J}\mathbf{v}$ is larger than $1$ for some $\mathbf{v}$ with $\|\mathbf{v}\|=1$, then the fixed point is unstable. For the rest of this Appendix, we evaluate the stability of Eqs. (3.1) with duplicated order parameters. Keeping $\mathbf{\bar{p}}$ and $\mathbf{p}$ fixed, the Jacobian of the saddle-point equation for $\mathbf{\bar{x}}$ is
@@
The subcovers $\mathcal{S}\left(0<\mu\leq P\right)$ and $\mathcal{S}\left(R<\mu\leq P+R\right)$ are both partitions of the indices $\mu_{*}$. As such, $\left[\partial_{\bar{x}^{\nu}_{k}}\bar{x}^{\mu}_{j}\right]_{\mu,\nu=1}^{P+R}$ is block diagonal with $2\times 2$ blocks coupling the indices $0<\mu\leq R$ and $\nu=\mu+P$. Without loss of generality, we investigate the stability of the block
@@
Appendix J Weights learned with unsupervised training
Appendix K Splitting steepest descent
@@
where $\bar{\nabla}_{\bm{\theta}}$ is the gradient constrained to the unit hypersphere $S^{N-1}$ (see Appendix C) and $\beta_{\text{eff}}$ can be written explicitly in terms of $\beta$ as $\beta_{\text{eff}}=\varsigma\beta$ (see Eqs. 12 and 3.1). At step 11 of Alg. (1), we break permutation symmetries in the memories $\mathbf{w}^{\mu}$ by descending along the eigenvectors $\mathbf{u}^{\mu}\in S^{N-1}$ corresponding to the eigenvalues $\lambda_{\text{min}}^{\mu}$. To be more precise, we update the memories $\mathbf{w}^{\mu}$ and their duplicates $\mathbf{w}^{\text{dupli},\mu}$ according to $\mathbf{w}^{\mu}\leftarrow\mathbf{w}^{\mu}+\delta\mathbf{u}^{\mu}$ and $\mathbf{w}^{\text{dupli},\mu}\leftarrow\mathbf{w}^{\text{dupli},\mu}-\delta\mathbf{u}^{\mu}$, respectively, where $\delta$ is a relatively small learning rate [29]. In physics terminology, the eigenvectors $\mathbf{u}^{\mu}$ are excitation modes that break the permutation symmetries of the memories. $N$ and $P$ are generally large, so it is prohibitively expensive to store $\mathcal{S}_{\mu}\left(\mathbf{w},\mathbf{p}\right)$ explicitly for all hidden units $1\leq\mu\leq P$. Therefore, as proposed in [75], we find the eigenvectors $\mathbf{u}^{\mu}$ and their eigenvalues $\lambda_{\text{min}}^{\mu}$ by minimizing the Rayleigh quotients $Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]:S^{N-1}\ni\mathbf{u}^{\mu}\mapsto\left[\mathbf{u}^{\mu}\right]^{T}\mathcal{S}_{\mu}\left(\mathbf{w},\mathbf{p}\right)\mathbf{u}^{\mu}$, which can be evaluated without constructing $\mathcal{S}_{\mu}\left(\mathbf{w},\mathbf{p}\right)$ explicitly. To derive $Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]$, we first compute $\bar{\nabla}_{\mathbf{w}^{\mu}}\bar{\nabla}_{\mathbf{w}^{\mu}}^{T}\exp\left(\beta_{\text{eff}}\sum_{i}w^{\mu}_{i}x_{i}\right)$. Given $f\left(\bm{\theta}\right)=\sum_{i}\theta_{i}x_{i}$, we find
@@
As mentioned at the beginning of Appendix C, $\bar{\nabla}_{\bm{\theta}}f\left(\bm{\theta}\right)$ is the unrestricted gradient of $f\left(\bm{\theta}\right)$ projected onto the tangent space of $\bm{\theta}$. It remains to calculate $\bar{\nabla}_{\bm{\theta}}\bar{\nabla}_{\bm{\theta}}^{T}f\left(\bm{\theta}\right)$. Following [89], we find it to be the iterated projected gradient
@@
This technique is based on the automatic differentiation trick proposed in [75]. We numerically implement the limit in the first equation by setting $\epsilon=0$ during loss evaluation and $\epsilon=1$ during gradient computation. When we optimize $\mathcal{Q}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}\right)$, the minima of the different Rayleigh quotients $Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)$ can span multiple orders of magnitude, so we normalize the gradients $\nabla_{\mathbf{u}^{\mu}}Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)$ by a running average of their magnitudes to facilitate convergence. This is inspired by the use of RMSProp in [75]. Moreover, we constrain the eigenvectors $\mathbf{u}^{\mu}$ and the gradients $\nabla_{\mathbf{u}^{\mu}}Q_{\mu}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)$ to the unit hypersphere as we do for $\mathbf{w}^{\mu}$ (see Appendix C). To make Fig. (7), we train the eigenvectors for $1$ epoch, during which we monitor $\min_{\mu}\left\{\frac{1}{2}\left[\mathbf{u}^{\mu}\right]^{T}\nabla_{\mathbf{u}^{\mu}}\mathcal{Q}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)\right\}=\min_{\mu}\left\{\mathcal{Q}\left[\mathbf{w},\mathbf{p}\right]\left(\mathbf{u}^{\mu}\right)\right\}\sim\min_{\mu}\left\{\lambda_{\text{min}}^{\mu}\right\}$ as a metric. The Rayleigh quotients converge very quickly, and further training of the eigenvectors is not necessary.